A collection of 34 coins consisting of dimes and quarters has a value of $5.50.
how many of each type of coin is there?
: A collection of 34 coins consisting of dimes and quarters has a value of $5.50.
how many of each type of coin is there?

Answer by jojo14344(879)

D+Q=34coins -------------------------> eqn 1
In eqn 1, we get D=34-Q -------------> eqn 2
Also,
0.10D+0.25Q=5.50 --------------------> eqn 3
Substitute eqn 2 in eqn 3:
0.10(34-Q)+0.25Q=5.50
3.4-0.10Q+0.25Q=5.50
0.15Q=5.50-3.40=2.10
cross(0.15)Q/cross(0.15)=cross(2.10)14/cross(0.15)
Q=14coins ---------> no. of Quarters
Go back eqn 2:
D=34-14=20 --------> no. of Dimes
in doubt? go back eqn 3:
0.10(20)+0.25(14)=5.50
2.00+3.50=5.50
5.50=5.50
Thank you,
Jojo